# Accelerating Ray Tracing-Based Wireless Channels Generation for Real-Time Network Digital Twins

### v1

@@ VII Conclusions and Future Works
This article posed two problems related to channel generation with RT and a methodology to assess their solutions. This methodology relies on constructing “ground truth” results, which allows evaluations using objective figures of merit such as NMSE. It was also described how the two problems fit modules that compose a NDT based on RT.
@@
This work also proposed different methods to address 3D scenario simplification (P2). Solutions to P2 are crucial to make feasible running current RT software for scenarios that represent large areas and/or have a large number of meshes. As solutions to P2, we proposed four different cut-out based methods that discard unnecessary regions, aiming that they do not impact the corresponding wireless channel. These methods explore the trade-off between the realism of the generated channel and the time to process a 3D scenario. Our cut-out based method results demonstrated that it is possible to decrease the level of unnecessary details and reduce the simulation by half, while retaining an accurate representation of the channel with respect to the ground truth.
Another contribution of this work is to release the source code of proposed methods and evaluation methodology. Upon acceptance, this software will be available at a Git repository. [^4] https://github.com/lasseufpa/speedup-rt
For future work, we will provide results obtained by combining solutions to both problems P1 and P2. When these solutions operate together in modules such as WCG and SOH, one can achieve almost real-time execution for relatively complex 3D scenarios in NDTs based on RT. We will also analyze the impact of the proposed methods on various communication metrics, particularly link-level indicators such as signal-to-noise ratio and channel capacity. Another ongoing work is the integration of the proposed methods into the CAVIAR framework [16]. Efficient RT pre and post-processing methods enable CAVIAR to execute system-level simulations using realistic site-specific wireless channels, and evaluations based on key performance indicators such as delay, jitter, and packet loss.
Acknowledgment
This study was financed in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq); the Innovation Center, Ericsson Telecomunicações Ltda., Brazil; Universal (CNPq grant 405111/2021-5) and Project Smart 5G Core and MUltiRAn Integration (SAMURAI) (MC-TIC/CGI.br/FAPESP under grant 2020/05127-2).

### v2

@@ VII Conclusions and Future Works
This article posed two problems related to channel generation with RT and a methodology to assess their solutions. This methodology relies on constructing “ground truth” results, which allows evaluations using objective figures of merit such as NMSE. It was also explained how the two problems map to modules in an enabler tool for wireless channel modeling in an RT-based NDT system. [^4] Source code will be available at: https://github.com/lasseufpa/speedup-rt
@@
This work also proposed different methods to address 3D scenario simplification (P2). Solutions to P2 are crucial to make feasible running current RT software for scenarios that represent large areas and/or have a large number of meshes. As solutions to P2, we proposed four different cut-out based methods that discard unnecessary regions, aiming that they do not impact the corresponding wireless channel. These methods explore the trade-off between the realism of the generated channel and the time to process a 3D scenario. Our cut-out based method results demonstrated that it is possible to decrease the level of unnecessary details and reduce the simulation time by half, while retaining an accurate representation of the channel with respect to the ground truth.
Finally, for future work, we will provide results obtained by combining solutions to both problems P1 and P2. When these solutions operate together in modules such as WCG and SOH, one can achieve almost real-time execution for relatively complex 3D scenarios in NDTs based on RT. We will also analyze the impact of the proposed methods on various communication metrics, particularly link-level indicators such as signal-to-noise ratio and channel capacity. Another ongoing work is the integration of the proposed methods into the CAVIAR framework [16]. Efficient RT pre and post-processing methods enable CAVIAR to execute system-level simulations using realistic site-specific wireless channels, and evaluations based on key performance indicators such as delay, jitter, and packet loss.